On January 1, 2017, Mcllroy, Inc., acquired a 60 percent interest in the common stock of Stinson, Inc., for $334,200. StinsonStinson, Inc. $ (357,000) 218, 200 74,400 0 $ (64,400) McIlroy, Inc. $ (720,000) 473,200 194,415 (33,071) $ (85,456) $ (761,7Complete this question by entering your answers in the tabs below. Required A Required B Show how McIlroy determined the $390NCI MCILROY, INC., AND STINSON, INC. Consolidation Worksheet For Year Ending December 31, 2018 Consolidation Entries Mcllroy$ $ $ $ $ (329,000)| 149,800 130,600 Retained earnings, 12/31 Cash and receivables Inventory Investment in Stinson Buildings
Answer #1

Part A

Determination of Investment in Stinson account balance

Consideration transferred

334200

Increase in Stinson's retained earnings

50520

Excess fair value amortization

(6594)

2017 ending inventory profit deferral

(10400)

33526

McIlroy's equity earnings in Stinson for 2018

33071

Stinson 2018 dividends declared to McIlroy

(10320)

Investment account balance 12/31/18

390477

(281800-197600)*60% = 38640

10990*60% = 6594

17200*60% =9900

Acquisition-date fair value allocation and excess amortizations

Consideration transferred

334200

Noncontrolling interest fair value

222800

Subsidiary fair value at acquisition-date

557000

Acquisition-date book value

297600

Fair value in excess of book value

259400

Life in Years

Annual Excess Amortizations

Excess fair value assignments

to patents

75700

10

7570

to customer list

51300

15

3420

to goodwill

132400

indefinite

-

10990

100000+197600 = 297600

259400-75700-51300 = 132400

2017

2018

Intra-entity profits (downstream)

52000

37500

Intra-entity transfers remaining in inventory

20%

25%

Gross profit rate

10400

9375

1-(124800/156000) = 20%

1-(112500/150000)= 25%

Stinson's 2018 income

64400

Excess fair value amortization

(10990)

Adjusted net income

53410

McIlroy's percentage ownership

60%

McIlroy's share of Stinson's adjusted net income

32046

2017 Intra-entity inventory profit recognized

10400

2018 Intra-entity inventory profit deferred

(9375)

McIlroy's equity earnings in Stinson

$33071

Part B

Consolidation Worksheet

Year Ending December 31, 2018

Adjustments & Eliminations

McIlroy

Stinson

Debit

Credit

Noncontrolling interest

Consolidate Totals

Sales

(720000)

(357000)

150000

(927000)

Cost of goods sold

473200

218200

9375

160400

540375

Operating expenses

194415

74400

10990

279805

Equity in earnings of Stinson

(33071)

0

33071

0

Separate company net income

(85456)

(64400)

Consolidated net income

(106820)

To noncontrolling interest

(21364)

21364

To McIlroy, Inc.

(85456)

Retained earnings, 1/1

(761700)

(281800)

281800

(761700)

Net income

(85456)

(64400)

(85456)

Dividends declared

47000

17200

9900

7300

47000

Retained earnings, 12/31

(800156)

(329000)

(800156)

Cash and receivables

272700

149800

422500

Inventory

256100

130600

9375

377325

Investment in Stinson

390477

0

9900

400377

0

Buildings (net)

332000

203900

535900

Equipment (net)

237100

87800

324900

Patents (net)

0

22300

75700

7570

90430

Customer list

51300

3420

47880

Goodwill

132400

132400

Total assets

1488377

594400

1931335

Liabilities

(388221)

(165400)

(553621)

Common stock

(300000)

(100000)

100000

(300000)

Noncontrolling interest 1/1

263494

263494

Noncontrolling interest 12/31

(277558)

Retained earnings 12/31

(800156)

(329000)

(800156)

Total liabilities and equity

(1488377)

(594400)

854536

854536

(1931335)